A right triangle has sides 15cm and 20cm. What is the lenght of the hypotenuse?
Answer by smoothflyin(6)

Pythagorean Theorum - (A^2+B^2=C^2)
-
C=Hypotenuse
15^2+20^2=625
Take the Square root of 625 and you have 25, the length of your Hypotenuse is 25cm, because
15^2+20^2=25^2